# Experiencer, Helper, or Observer: Online Fraud Intervention for Older Adults Through Role-based Simulation

### v1

@@ 4.4.2. User Experience Evaluation
After finishing the online fraud awareness assessment, participants were asked to report their learning experience with the assigned intervention. This evaluation was conducted through a survey with details shown in the Appendix B. The survey was implemented on our application following the awareness tests, and participants’ responses were automatically logged.
@@ 4.5. Follow-up Intervention Evaluation
We reconnected with the collaborating community center, which assisted us in inviting participants who had taken part in the main study to voluntarily return for a follow-up session approximately one month later. [^6] The timing was around one month rather than an exact day, as the main study itself spanned several weeks, and we could not ensure that each participant returned on precisely the same day. A total of 89 participants completed the follow-up study, of which 5 were excluded due to repeated tests, resulting in 84 valid users, including 17 participants in the Experiencer group, 31 participants in the Helper group, 17 participants in the Observer group, and 19 participants in the Control group. The demographic information of these older adults is shown in the Supplementary Material.
@@ 5.1. Quantitative Data Analysis
We first examined data completeness and excluded cases with missing values. Randomization of group assignment was then assessed across demographic factors, including age, gender, and education level. For age, normality (Shapiro–Wilk test) and homogeneity of variance (Levene’s test) were evaluated. As the assumptions for ANOVA were not met, the non-parametric Kruskal–Wallis test was applied. Gender distribution was examined using a chi-square test of independence. Education level was assessed using Fisher’s exact test with Monte Carlo simulation due to sparse cells (expected counts smaller than five), which provides a robust alternative to the chi-square test for larger contingency tables.
@@
To further examine group-level differences, we conducted two sets of Kruskal–Wallis H tests (Kruskal and Wallis, 1952). First, we compared outcomes across all four groups (Experiencer, Helper, Observer, Control) to identify differences between the treatment and control groups. Second, we compared outcomes across the three treatment groups (Experiencer, Helper, Observer) to investigate potential role-specific effects. This non-parametric test is appropriate for independent groups when the dependent variable is ordinal or continuous but not normally distributed. When the Kruskal–Wallis test indicated significance, we conducted post-hoc pairwise comparisons using Dunn’s test with Bonferroni correction to adjust for multiple testing. All quantitative data analyses and visualizations were performed in R and Python.
5.1.4. Follow-up Analysis.
Main vs. follow-up (within-group changes). To examine whether the effectiveness of the intervention was retained over time in each group, we compared participants’ fraud cue identification and self-efficacy between the main study and the follow-up. When data were approximately normally distributed, we used paired $t$-tests; otherwise, we applied Wilcoxon signed-rank tests.
Follow-up (between-group differences). We applied the same methods to compare group differences in the follow-up evaluation as in the main study, shown in Sec. 5.1.3.
@@ 6.1. Preliminary Validation Analysis
The results of a CFA indicated an acceptable model fit ($\chi^{2}$/df = 2.06, CFI=0.915, TLI=0.894, RMSEA=0.086, SRMR=0.072). All constructs exhibited Cronbach’s alpha values near or above .70, demonstrating acceptable reliability. Two constructs, usability and self-efficacy, had slightly lower values of 0.634 and 0.666, respectively. While 0.7 is often cited as the benchmark for acceptable internal consistency, recent methodological reviews note that values approaching this threshold can still be considered adequate, particularly in exploratory or applied contexts (Taber, 2018). Given that both values are close to 0.7, we retained these constructs in the analysis.
@@ 6.2. Effectiveness of Intervention
As shown in Fig. 5(a), participants in all treatment groups (Experiencer, Helper, Observer) achieved higher average $F_{1}$ scores in the online fraud cue recognition compared to the control group. A Mann–Whitney U test confirmed that these differences were statistically significant (W=1009.5, P <0.001). A Kruskal–Wallis test revealed significant differences across the four groups (H(3)=24.68, P <0.001), shown in Tab. 2. Post-hoc pairwise comparisons using Dunn’s test with Bonferroni correction identified a significant difference between the Helper and Control group (Z=4.73, adjusted P <.001), as well as between the Experiencer and Control group (Z=3.53, adjusted P=.002), displayed in Tab. 3.
@@
treatment vs. control fraud cue identification 24.68 <.001***
@@
among treatment fraud cue identification 6.97 .031*
@@
Among the three treatment conditions, the Helper group achieved the highest mean $F_{1}$ score, followed by the Experiencer group, while the Observer group had the lowest in Figure 5(a). The Kruskal–Wallis test indicated a significant difference among the three treatment groups (H(2)=6.97, P=.031) in Tab. 2. Dunn’s post-hoc tests with Bonferroni adjustment showed that Helpers outperformed Observers (Z=2.63, adjusted P=.025), whereas no significant differences were observed between Helpers and Experiencers or between Experiencers and Observers, shown in Tab. 4.
@@
All three treatment groups showed higher mean self-efficacy scores than the control group as shown in Fig. 5(b). Nonetheless, the differences were not statistically significant, as indicated by both the Mann–Whitney U test (W=1737.5, P=.314) and the Kruskal–Wallis test (H(3)=1.98, P=.576) in Tab. 2. The mean self-efficacy scores were 4.41 for Experiencer, 4.40 for Helper, and 4.26 for Observer in Fig. 5(b). However, the Kruskal–Wallis test revealed no significant differences across these groups (H(2)=1.09, P=.581) in Tab. 2.
@@ 6.3. User Experience of Intervention
Fig. 6 presents the descriptive results for user experience measures. Overall, the Mann–Whitney U tests revealed no statistically significant differences between the treatment groups and the control group across any of the constructs. Likewise, the Kruskal–Wallis tests indicated no significant differences among the four groups. Tab. 5 reports the detailed Kruskal–Wallis results for user experience.
@@
The Kruskal–Wallis test revealed no significant differences among Experiencer, Helper, and Observer groups for any user experience construct, shown in Tab. 5. Although the Observer group showed the highest means for Usability and Future Behavioral Intention, neither of these differences reached statistical significance.
@@
6.4. One-month Follow-up Effects of Intervention
Fig. 7 demonstrates that scores in all training conditions remained relatively stable from the main study to the follow-up, while the control group exhibited a modest, non-significant increase. Moreover, for fraud cue identification, no significant within-group differences were found between the main and follow-up assessments across all groups, including Control (Z=1.660, P=.102), Experiencer (Z=-0.284, P=.798), Helper (t=-0.582, P=.565), and Observer (t=0.331, P=.745). Similarly, for self-efficacy, no significant changes were observed over time within any group, including Control (Z=-0.118, P=.937), Experiencer (Z=-0.280, P=.831), Helper (Z=0.560, P=.582), and Observer (t=-0.347, P=.733), shown in Tab. 6.
(a) Fraud cue identification (mean ± SD)
(b) Self-efficacy (mean ± SD)
Group Shapiro p-value Test-used Test Statistic P-value
helper 0.412 Paired t-test t = -0.582 .565
observer 0.689 Paired t-test t = 0.331 .745
control 0.023 Wilcoxon signed-rank Z = 1.660 .102
helper 0.010 Wilcoxon signed-rank Z = 0.560 .582
observer 0.102 Paired t-test t = -0.347 .733
control 0.023 Wilcoxon signed-rank Z = -0.118 .937
Mann-Whitney U tests did not find a significant difference for both fraud cue identification (W=489, P=.169) and self-efficacy (W=602, P=.863) between treatment groups and the control group. Similarly, Kruskal-Wallis tests indicated no significant difference, with specific results in Tab. 7. Moreover, there is no significant effectiveness difference between the three treatment groups in Tab. 7.
Effectiveness Construct Test Statistic P-value
treatment vs. control fraud cue identification 2.95 .399
self-efficacy 2.07 .558
among treatment fraud cue identification 0.98 .613
self-efficacy 2.02 .364
6.5. User Feedback of Intervention
6.5.1. Perceived Benefits of LLM-based Simulation
This approach to fraud prevention feels more detailed, more targeted, and also more accurate, with a lot of concrete details. In all these details, you can see how to resist such behaviors. Today’s (LLM-based simulation) was quite detailed. When I give different content and prompts, its responses also vary. That way of chatting makes the experience sink in more deeply. It really sticks in the mind and leaves a stronger impression.
Yours is definitely more flexible. Compared to traditional lectures, this one covers a much broader range. Lectures are really limited, narrow, and rigid, at least, very rigid. But with your system, I can let my thoughts wander, and it can also respond in unexpected ways, right? It’s more like that, because when you ask, it will always answer, isn’t that so? In that sense, the flexibility and liveliness are definitely much better.
For someone like me, who is hard to trick, your system didn’t force me to keep chatting. With real scammers, the ones actually trying to get money, it would feel more exhausting. With the AI, I could just say a few words and then stop. I know very clearly that if something is a scam, I normally wouldn’t believe it. So, I politely refused. The AI doesn’t mind and won’t keep chasing you, and you can still learn some knowledge about scams from it.
Finally, several participants envisioned future uses of AI in decision support. They imagined shifting from the helper role, where older adults assist a potential victim, to AI directly supporting their own decision-making. For example, P109 suggested, I think it (AI) is definitely more skilled than me in giving action advice. What if the AI persuades me instead? When facing suspicious situations, I could use AI to help me solve them.” P118 similarly hoped that if a friend were scammed, she could use AI to help persuade them more effectively, “If I realize my friends might be getting scammed and can’t convince you, I could bring in a robot to help persuade you. Because when you call the police, sometimes they can’t give that many reasons. But a robot could be added to the conversation to help persuade your friend.”
6.5.2. Perceived Limitations of LLM-based Simulation.
It’s like a system, a programmed system, because it lacks the warmth that a real person has. You know, when scammers try to deceive, they put warmth into their language. So, I feel that in terms of language, that aspect is stronger than what you have here.
6.5.3. Experiencer Feedback
Why go and learn from them? All they do is set traps for you. The more you chat with them, the more they start to use tricks. The deeper they go with their tricks, the more you keep chatting, and in the end, you’ll naturally start to believe them and get carried away.
P18 said, “They are scammers, deceivers. Asking me to chat with them, but I don’t even want to look at it.” However, when we emphasized the educational purpose again, the same participant (P18) acknowledged, “If it is about reminders and warnings, then I can fully accept it.”
6.5.4. Helper Feedback
Helping others gave me a sense of pleasure and altruism. When it comes to preventing fraud, if older adults are being deceived, and you have good experience to share, why not provide it, as long as it’s the right way? If you can reduce the chance of someone around you being scammed, why wouldn’t you do it? You don’t suffer any loss, right? And besides, in China we advocate for helping others with joy.
Since you asked me to persuade him, I just told him it was a scam, a fraud. But let’s say he really, really believed in it, and he kept circling around that topic and wouldn’t let go. With people, after you say a couple of sentences like ‘this is a scam’ a normal person might stop and think, ‘Oh, maybe I should reconsider.’ But here, it felt like he was too deeply poisoned. It’s not the same as chatting with a real person. Because with the system, one question will just keep coming back again and again.
6.5.5. Observer Feedback
Others were scammed. You need to reflect and summarize, right? Sometimes if you don’t reflect, you won’t notice, and you’ll get scammed yourself. You have to draw lessons from others’ experiences. Honestly, no matter how smart you are, if you’re not paying attention, you can still be scammed.
7. Discussion
7.1.3. Generalizability
Our fraud cues are derived from real scam cases and grounded in the specific vulnerabilities of older adults. For example, the cue “high returns with low risk” corresponds to the vulnerability related to financial expectations. Many different fraud types are in fact built upon the same underlying vulnerabilities (Deng et al., 2025), which means that the cues we identified can be, to some extent, generalized to educational interventions targeting other fraud types that exploit financial expectations or health and well-being concerns in real-world scenarios. More broadly, our work offers a transferable approach that extracting fraud cues through the lens of user vulnerabilities and using these cues as the basis for fraud education. This approach could be adopted by future educators or researchers who aim to design fraud-prevention interventions, especially in chat-based, step-by-step manipulative scam contexts. They could apply the same logic to identify fraud cues associated with the specific vulnerability they intend to target, and then use those cues to support tailored educational goals. Moreover, because the fraud cues and learning materials in our study were based on the Chinese context, the extent to which role-based simulation for online fraud education can be generalized to other cultural settings remains an open question. Future work could build on fraud cues and prevalent fraud contexts in other regions to examine the effectiveness of such interventions across cultures. Beyond cultural adaptation, generalizability to other populations is also worth exploring. Prior research has suggested that younger populations tend to be receptive to chatbot-based and interactive learning (Aly et al., 2024), and simulated phishing has already been studied with general populations (Nasser et al., 2020; Downs et al., 2006; Sheng et al., 2007). Thus, LLM-based role-based simulation may also hold promise beyond older adults.
7.2. Broadening Fraud Education through Multi-Role Perspectives
7.2.1. Experiencer
7.2.2. Helper
Despite the benefits of the helper role, several challenges deserve attention in future work. Some helpers expressed doubts about their role, including concerns that the person they were assisting might actually be a scammer attempting to entrap them (see Sec. 6.5.4). This highlights the importance of clearly communicating to helpers that the other party is a victim in need of assistance, rather than a potential threat. One potential solution is to particularly note differences between task-related interactions and prior background information. Besides, some helpers reported disappointment when their persuasion efforts were unsuccessful (Sec. 6.5.4). Future work should pay careful attention to the emotional burden that the helper role may place on older adults, particularly when they encounter scenarios involving distressed victims or feel responsible for others’ safety. In order to preserve older adults’ confidence, future iterations may benefit from a gamified design, where persuasion tasks are structured with varying levels of difficulty to ensure that participants experience achievable successes while still being challenged.
7.2.3. Observer
Older adults reported that simply observing produced a warning effect and enabled empathy and self-reflection (Sec. 6.5.5). This aligns with the principles of social learning theory, which posits that observation alone without direct practice can still facilitate learning (Bandura and Walters, 1977). However, our quantitative results did not show that observers outperformed participants in other conditions. One possible reason is that, without interactive engagement, older adults may only gain a general sense of the risks of online fraud while struggling to pinpoint specific cues of danger. Prior work suggests that observation can be more effective when supported by appropriate observing tools (O’Regan et al., 2016). Building on this, future designs could provide observers with lightweight scaffolds, for instance, a visual annotation tool or mini-quizzes alongside the interaction, that might help them more actively process and retain what they observe.
7.3. Implications
7.3.1. Design Implications
Integrating different roles into one application. Preferences for different roles may not only stem from the roles themselves but also from individual characteristics such as personality. For example, we observed that some older adults in the experiencer group reported that they were more introverted and therefore did not enjoy chat-based interactions (Sec. 6.5.3). For them, the observer role might have been more suitable. Since our study followed a between-group design, each participant engaged with only one role, and thus we could not directly compare within-person preferences across roles. Future systems could integrate multiple roles into a single educational application, allowing older adults to freely select and experiment with different roles. Such flexibility could help individuals discover one or more formats that best align with their preferences and allow them to approach the same problem from multiple perspectives to support more well-rounded learning.
Scaling role-based interventions. Future work could explore the scalability of the current intervention. Our experimental design focused on two scenarios, health and finance, but many other fraud contexts exploit older adults’ vulnerabilities, such as overreliance on authority, social isolation, and generational affection (see Sec. 3.2.3). These scenarios could be readily incorporated into the intervention by organizing additional fraud cues. Moreover, the scalability could also extend to the simulation channels. In our current design, WeChat served as the messaging interface for role-based simulation. However, livestreaming platforms represent another major site of online fraud targeting older adults. Such platforms could similarly support multi-role perspectives, for instance, by simulating audiences or content moderators as additional roles. Finally, scalability might also apply to the interaction patterns within each role. For example, in the helper condition, “learning by teaching” can take multiple forms, including teaching one peer, teaching a group of peers, or teaching collaboratively in a group where each member becomes an expert on a different topic (Slavin, 1980; Debbané et al., 2023). Future research could explore whether adapting these different forms, such as helping one victim, assisting a group, or collaborating with other older adults to provide help, might further improve learning outcomes in online fraud education.
7.3.2. Educational Implications
Advocating for the longitudinal effects. Role-based simulation learning proved more effective than the control condition after the intervention (see Sec. 6.2.1). However, the follow-up evaluation did not reproduce the significant group differences found in the main study (see Sec. 6.4). Although the follow-up results diverged from the initial pattern, this finding underscores the importance of incorporating longitudinal evaluation when studying online fraud education. Cross-sectional post-intervention gains may not fully reflect the stability or durability of learning outcomes, and without follow-up assessments, potential fluctuations, whether due to sampling constraints or other contextual factors, may go unnoticed. Our reflection extends to prior educational intervention studies, where evaluations of longitudinal effects are often lacking. Therefore, we advocate for future research to systematically incorporate longitudinal evaluation in order to capture the longitudinal effects of learning outcomes and better inform the design of durable and effective intervention strategies.
7.4. Limitations
8. Conclusion
Appendix A Fraud Cue Identification Test
A.1. Fraud Cue Identification Test (Main Study)
Finance A Test
Figure 8. Finance A Test
  • Promise high returns with low risk
  • Claim to be financial experts
  • Leverage peer influence
  • Pitch senior-living investment schemes
  • Invoke fake government investment schemes
  • Promise dividends between shareholders
  • None of the above
Finance B Test
Figure 9. Finance B Test
  • Promise high returns with low risk
  • Claim to be financial experts
  • Leverage peer influence
  • Pitch senior-living investment schemes
  • Invoke fake government investment schemes
  • Promise dividends between shareholders
  • None of the above
Health A Test
Figure 10. Health A Test
  • Exaggerating illness to create anxiety
  • Advertise referrals to top doctors
  • Offer free services as bait
  • Using cheap trials with steep follow-up charges
  • Claim to be medical experts
  • Framing products as miracle cures with refund guarantees
  • None of the above
Health B Test
Figure 11. Health B Test
  • Exaggerating illness to create anxiety
  • Advertise referrals to top doctors
  • Offer free services as bait
  • Using cheap trials with steep follow-up charges
  • Claim to be medical experts
  • Framing products as miracle cures with refund guarantees
  • None of the above
A.2. Fraud Cue Identification Test (Follow-up Study)
Finance A Test
Figure 12. Finance A Test
  • Promise high returns with low risk
  • Claim to be financial experts
  • Leverage peer influence
  • Pitch senior-living investment schemes
  • Invoke fake government investment schemes
  • Promise dividends between shareholders
  • None of the above
Finance B Test
Figure 13. Finance B Test
  • Promise high returns with low risk
  • Claim to be financial experts
  • Leverage peer influence
  • Pitch senior-living investment schemes
  • Invoke fake government investment schemes
  • Promise dividends between shareholders
  • None of the above
Health A Test
Figure 14. Health A Test
  • Exaggerating illness to create anxiety
  • Advertise referrals to top doctors
  • Offer free services as bait
  • Using cheap trials with steep follow-up charges
  • Claim to be medical experts
  • Framing products as miracle cures with refund guarantees
  • None of the above
Health B Test
Figure 15. Health B Test
  • Exaggerating illness to create anxiety
  • Advertise referrals to top doctors
  • Offer free services as bait
  • Using cheap trials with steep follow-up charges
  • Claim to be medical experts
  • Framing products as miracle cures with refund guarantees
  • None of the above
Appendix B Scale Questions in Evaluation
Table 8. Measurement items used in the evaluation.
Construct Items
self-efficacy [81] I am confident I can detect online scams.
I am able to detect online scams without much effort.
enjoyment [2] I have fun interacting with the system.
Using the system provides me with a lot of enjoyment.
I enjoy using the system.
Using the system bores me.
usability [41] This learning system’s capabilities meet my requirements.
Using this learning system is a frustrating experience.
This learning system is easy to use.
future behavioral intention [2] I plan to use the system in the future.
I intend to continue using the system in the future.
I expect my use of the system to continue in the future.
Appendix C User-System Interaction
In this section, we describe the interaction statistics between users and ROLESafe. Tab. 9 demonstrates the cue coverage rates across groups and contexts during the learning phase. For the experiencer and helper groups, since they interact with the LLM directly, the entire process remains dynamic, so we manually annotated the cue coverage. Specifically, if the LLM produced an utterance that included a particular fraud cue, we considered that conversation to have covered that cue. Conversely, if none of the LLM’s responses contained the cue, the conversation was coded as not covering it. For example, in Fig. 3 (a), the LLM’s statement “this is a government-supported program” is counted as covering the fraud cue of “invoke fake government investment schemes” within that conversation. To establish a shared coding standard, two authors annotated the presence or absence of covered cues. To assess inter-rater reliability, the two authors independently coded a random sample of 70 conversations. The results showed a Cohen’s Kappa of at least 0.9 for every cue, indicating strong agreement between coders. Afterward, the coders discussed discrepancies and reached consensus on the annotations. Each author then individually coded half of the remaining conversations. For the observer group, participants read pre-generated fraud cases rather than interacting with the LLM, so each fraud case includes all three cues.

### v2

@@ 4.4.2. User Experience Evaluation
After finishing the online fraud awareness assessment, participants were asked to report their learning experience with the assigned intervention. This evaluation was conducted through a survey (details in Appendix B). The survey was implemented on our application following the awareness tests, and participants’ responses were automatically logged.
@@ 4.5. Follow-up Intervention Evaluation
We reconnected with the collaborating community center, which assisted us in inviting participants who had taken part in the main study to voluntarily return for a follow-up session approximately one month later. [^6] The timing was around one month rather than an exact day, as the main study itself spanned several weeks, and we could not ensure that each participant returned on precisely the same day. A total of 89 participants completed the follow-up study, of which five were excluded due to repeated tests, resulting in 84 valid users, including 17 participants in the Experiencer group, 31 participants in the Helper group, 17 participants in the Observer group, and 19 participants in the Control group. The demographics of these older adults is shown in the Supplementary Material.
@@ 5.1. Quantitative Data Analysis
We first examined data completeness and excluded cases with missing values. Randomization of group assignment was then assessed across demographic factors, including age, gender, and education level. For age, normality (Shapiro–Wilk test) and homogeneity of variance (Levene’s test) were evaluated. As the assumptions for ANOVA were not met, the non-parametric Kruskal–Wallis test was applied. Gender distribution was examined using a chi-square test of independence. Education level was assessed using Fisher’s exact test with Monte Carlo simulation due to sparse cells (expected counts smaller than five) as an alternative to the chi-square test for larger contingency tables.
@@
To further examine group-level differences, we conducted two sets of Kruskal–Wallis H tests (Kruskal and Wallis, 1952). First, we compared outcomes across all four groups (Experiencer, Helper, Observer, Control) to identify differences between the treatment and control groups. Second, we compared outcomes across the three treatment groups (Experiencer, Helper, Observer) to investigate any role-specific effects. This non-parametric test is appropriate for independent groups when the dependent variable is ordinal or continuous but not normally distributed. When the Kruskal–Wallis test indicated significance, we conducted post-hoc pairwise comparisons using Dunn’s test with Bonferroni correction to adjust for multiple testing.
5.1.4. Follow-up Study Analysis.
To examine whether the effectiveness of the intervention was retained over time in each group, we compared participants’ fraud cue identification and self-efficacy between the main study and the follow-up. When data were approximately normally distributed, we used paired $t$-tests; otherwise, we applied Wilcoxon signed-rank tests. We applied the same methods to compare group differences in the follow-up evaluation as in the main study, shown in Sec. 5.1.3.
@@ 6.1. Preliminary Validation Analysis
The results of a CFA indicated an acceptable model fit ($\chi^{2}$/df=2.06, CFI=0.915, TLI=0.894, RMSEA=0.086, SRMR=0.072). All constructs exhibited Cronbach’s alpha values near or above 0.7, demonstrating acceptable reliability. Two constructs, usability and self-efficacy, had slightly lower values of 0.634 and 0.666, respectively. While 0.7 is often cited as the benchmark for acceptable internal consistency, recent methodological reviews note that values approaching this threshold can still be considered adequate, particularly in exploratory or applied contexts (Taber, 2018). Given that both values are close to 0.7, we retained these constructs in the analysis.
@@ 6.2. Effectiveness of Intervention
As shown in 5(a), participants in all treatment groups (Experiencer, Helper, Observer) achieved higher average $F_{1}$ scores in the online fraud cue recognition compared to the control group. A Mann–Whitney U test confirmed that these differences were statistically significant (W=1009.5, P<0.001). A Kruskal–Wallis test revealed significant differences across the four groups (H(3)=24.68, P<0.001), shown in Table 2. Post-hoc pairwise comparisons using Dunn’s test with Bonferroni correction identified a significant difference between the Helper and Control group (Z=4.73, adjusted P<.001), as well as between the Experiencer and Control group (Z=3.53, adjusted P=.002), displayed in Table 3.
@@
Figure 5: Bar charts showing descriptive comparisons across groups. (a) Fraud cue identification (F1 scores) was higher in the treatment groups (Experiencer 0.81, Helper 0.85, Observer 0.76) than in the Control group (0.66). (b) Self-efficacy scores were similar across groups, with Experiencer (4.41), Helper (4.40), Observer (4.26), and Control (4.04).
@@
Table 2: Results of Kruskal–Wallis tests on intervention effectiveness. Compared to the control group, treatment groups showed significantly higher fraud cue identification (p \$<\$ .001), but no significant difference in self-efficacy. Among treatment groups, fraud cue identification also differed significantly (p = .031), while self-efficacy showed no significant difference.
@@
treatment vs. control fraud cue identification 24.68 $001^{***}$
@@
among treatment fraud cue identification 6.97 $031^{*}$
@@
Table 3: Pairwise comparisons of fraud cue identification across groups using Dunn’s test with Bonferroni correction. Both the Helper group (p \$<\$ .001) and Experiencer group (p = .002) performed significantly better than the Control group.
@@
Among the three treatment conditions, the Helper group achieved the highest mean $F_{1}$ score, followed by the Experiencer group, while the Observer group had the lowest in 5(a). The Kruskal–Wallis test indicated a significant difference among the three treatment groups (H(2)=6.97, P=.031) in Table 2. Dunn’s post-hoc tests with Bonferroni adjustment showed that Helpers outperformed Observers (Z=2.63, adjusted P=.025), whereas no significant differences were observed between Helpers and Experiencers or between Experiencers and Observers, shown in Table 4.
@@
Table 4: Pairwise comparisons among treatment groups on fraud cue identification using Dunn’s test with Bonferroni correction. The Helper group significantly outperformed the Observer group (p = .025). No significant differences were found between Experiencer and Helper or between Observer and Experiencer.
@@
All three treatment groups showed higher mean self-efficacy scores than the control group, as shown in 5(b). Nonetheless, the differences were not statistically significant, as indicated by both the Mann–Whitney U test (W=1737.5, P=.314) and the Kruskal–Wallis test (H(3)=1.98, P=.576) in Table 2. The mean self-efficacy scores were 4.41 for Experiencer, 4.40 for Helper, and 4.26 for Observer in 5(b). However, the Kruskal–Wallis test revealed no significant differences across these groups (H(2)=1.09, P=.581) in Table 2.
@@ 6.3. User Experience of Intervention
Figure 6 presents the descriptive results for user experience measures. Overall, the Mann–Whitney U tests revealed no statistically significant differences between the treatment groups and the control group across any of the constructs. Likewise, the Kruskal–Wallis tests indicated no significant differences among the four groups (see Table 5). Although the Observer group showed the highest means for Usability and Future Behavioral Intention, neither of these differences reached statistical significance.
@@
Figure 6: Bar charts comparing user experience across groups. (a) Enjoyment scores were similar across groups, ranging from 4.19 (Experiencer) to 4.48 (Control). (b) Usability ratings were also comparable, with means from 4.34 (Control) to 4.56 (Observer). (c) Future behavioral intention scores ranged from 4.08 (Helper) to 4.63 (Observer), with no large differences between groups.
@@
Table 5: Results of Kruskal–Wallis tests on user experience constructs. No significant differences were found between treatment and control groups or among treatment groups for enjoyment, usability, or future behavioral intention.
@@
6.4. User-System Interaction
In this section, we describe the interaction statistics between users and ROLESafe. Table 6 demonstrates the cue coverage rates across groups and contexts during the learning phase. For the experiencer and helper groups, since they interact with the LLM directly, the entire process remains dynamic, so we manually annotated the cue coverage. Specifically, if the LLM produced an utterance that included a particular fraud cue, we considered that conversation to have covered that cue. Conversely, if none of the LLM’s responses contained the cue, the conversation was coded as not covering it. For example, in Figure 3a, the LLM’s statement “this is a government-supported program” is counted as covering the fraud cue of “invoke fake government investment schemes” within that conversation. To establish a shared coding standard, two authors annotated the presence or absence of covered cues. To assess inter-rater reliability, the two authors independently coded a random sample of 70 conversations. The results showed a Cohen’s Kappa of at least 0.9 for every cue, indicating strong agreement between coders. Afterward, the coders discussed discrepancies and reached consensus on the annotations. Each author then individually coded half of the remaining conversations. For the observer group, participants read pre-generated fraud cases rather than interacting with the LLM, so each fraud case includes all three cues.